OpenStudy (anonymous):

\[\left| x+2 \right| + x + \left| x-1 \right|-3=0\]

OpenStudy (anonymous):

consider intervals , for x<-2 we have (-x-2) + x +(1-x) - 3 =0 for -2<x<1 we have x + 2 + x + (1-x) -3 = 0 for 1>x we have x+2 + x + x -1 -3 = 0

OpenStudy (anonymous):

so wat the interval period....

OpenStudy (anonymous):

basically consider each interval where the modulus functions change, and solve each equation for x<-2 we have (-x-2) + x +(1-x) - 3 =0 -x -2 + x + 1 -x -3 = 0 -x - 4 = 0 x = -4 now check if this works, -4<-2 so it is valid next you need to consider -2<x<1

OpenStudy (anonymous):

ok let me try and will inform u!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

thxx lol @eigenschmeigen i understod the method!

OpenStudy (anonymous):

@eigenschmeigen help http://openstudy.com/study?signup#/updates/4fd0936ee4b0e04663181e2a